Question: If Coke and Pepsi are perfect substitutes for Lynn. She is always willing to substitute 1.1 Pepsis for 1 Coke. That is, her utility function is: U(Cokes, Pepsis)= 1.1*Cokes + Pepsis

If she has an income of $100 and the price of Cokes is $1.20 and Pepsis is $1, then what is the optimal bundle that Lynn should consume?
Answer #1

MRS = MUc/ MUp = 1.1/1 = 1.1

Pc/Pp = 1.20/1 = 1.2

Because MRS<Pc/Pp . So, the optimal bundle is Lynn to spend all of her income on Pepsi only and not consume Coke.

Income = $100

So, the optimal bundle is (I/Pp) = 100/1 = 100 units.

Hence, optimal bundle is (C, P) = (0,100).